### v1

# HiCL: Hippocampal-Inspired Continual Learning

Abstract
We propose HiCL, a novel hippocampal-inspired dual-memory continual learning architecture designed to mitigate catastrophic forgetting by using elements inspired by the hippocampal circuitry. Our system encodes inputs through a grid-cell-like layer, followed by sparse pattern separation using a dentate gyrus-inspired module with top-$k$ sparsity. Episodic memory traces are maintained in a CA3-like autoassociative memory. Task-specific processing is dynamically managed via a DG-gated mixture-of-experts mechanism, wherein inputs are routed to experts based on cosine similarity between their normalized sparse DG representations and learned task-specific DG prototypes computed through online exponential moving averages. This biologically grounded yet mathematically principled gating strategy enables differentiable, scalable task-routing without relying on a separate gating network, and enhances the model’s adaptability and efficiency in learning multiple sequential tasks. Cortical outputs are consolidated using Elastic Weight Consolidation weighted by inter-task similarity. Crucially, we incorporate prioritized replay of stored patterns to reinforce essential past experiences. Evaluations on standard continual learning benchmarks demonstrate the effectiveness of our architecture in reducing task interference, achieving near state-of-the-art results in continual learning tasks at lower computational costs.
Introduction
Artificial neural networks excel at many single‐task benchmarks but struggle when trained sequentially on multiple tasks, suffering from catastrophic forgetting—the tendency to overwrite previous knowledge with new information. In contrast, the human brain learns continuously throughout life, preserving decades of memories while acquiring new skills.
In continual learning, every new task is like a distinct episode in our life’s unfolding story. Just as the hippocampus rapidly encodes unique experiences into episodic memory, tagging them with contextual and temporal cues, a continual learner must capture each task’s defining features before they fade. The brain then replays these episodes—during sleep or quiet wakefulness—to gradually integrate them into neocortical networks, preventing older memories from being overwritten by new ones. We leverage this complementary learning systems principle by treating each task as an episodic trace: we store compact representations in a replay buffer and periodically revisit them during training, mirroring how hippocampal replay consolidates previous and precious episodes into lasting knowledge [1].
Figure 1: Schematic of the HiCl architecture overlaid with the hippocampal trisynaptic circuit. The entorhinal cortex (layers II/III) projects to the dentate gyrus (DG), which performs pattern separation. DG outputs feed into CA3, an autoassociative network responsible for pattern completion. CA3 then projects to CA1, which integrates inputs from both CA3 and the entorhinal cortex before sending output to other brain regions. This circuit underlies memory encoding and retrieval processes. The general background here of the hippocampal circuit is based on [2] and adapted from [3, 4, 5, 6].
We draw on three core hippocampal subregions—the dentate gyrus (DG), CA3, and CA1—to inspire a DG‑Gated Mixture‑of‑Experts (MoE) architecture, where each subcircuit is instantiated as an expert and inputs are dynamically routed based on their DG activations. This DG‑based gating, eliminates the need for task labels at inference and directly tackles catastrophic forgetting.
Dentate Gyrus (DG): In biology, the DG performs pattern separation by mapping inputs onto a very sparse, typically only $3–5\%$ of granule cells fire for any given pattern [7]. In our model, the SparseActivation DG layer enforces top‑$k$ sparsity $(k=5\%)$, orthogonalizing features both for expert representations and for routing signals.
CA3: The DG’s orthogonal codes feed into CA3’s recurrent attractor network, which functions as pattern completion, reconstructing full memory traces from partial cues [8]. We mirror this with a lightweight two-layer MLP that refines and transforms DG outputs via a non-linear projection, functionally completing the pattern for downstream processing.
CA1: Finally, CA1 integrates completed patterns from CA3 with direct cortical inputs, acting as a comparator and coordinating memory consolidation through hippocampo‑cortical dialogue [9]. We emulate this via our ca1_integration block and a consolidation stage combining Elastic Weight Consolidation (EWC) with a replay buffer.
Bringing it together: By instantiating $N$ such experts and routing via cosine similarities between live DG codes and stored DG prototypes—using either hard or soft‑max gating—and then sharpening DG separation with a contrastive tuning phase.
This biology‑inspired solution to the stability–plasticity dilemma offers actionable insights for continual learning in artificial systems. Our DG‑Gated Mixture‑of‑Experts architecture translates hippocampal principles into three core mechanisms: (1) a top-$k$ sparse DG layer ($k\approx 5\%$) that orthogonalizes inputs for both expert encoding and dynamic routing; (2) a CA3‑inspired two‑layer MLP autoencoder that completes partial patterns via attractor-like reconstruction; and (3) a consolidation module combining a prioritized replay buffer with Elastic Weight Consolidation to protect prior knowledge.
By instantiating the hippocampal tripartite circuit across multiple experts and leveraging DG‑based gating to minimize interference, CA3‑style pattern completion for robust recall, and replay+EWC for selective consolidation, our DG‑Gated MoE exemplifies how neuroscience‑informed architectures can achieve efficient, adaptive continual learning.
Figure 2: Overview of the HiCL architecture aligned with the hippocampal trisynaptic circuit: EC II/III → DG (pattern separation) → CA3 (pattern completion) → CA1 (integration).
Our primary contributions are:
  • We introduce HiCL, a DG-gated Mixture-of-Experts (MoE) model that routes inputs without a gating network. By leveraging the sparse representations from a Dentate Gyrus-inspired layer, our model dynamically assigns tasks to experts based on cosine similarity to learned prototypes.
  • We developed a training strategy that explicitly improves pattern separation. Our two-phase approach first trains individual experts using EWC, replay, and feature distillation, then applies a global contrastive loss to all DG layers, forcing them to learn more distinct representations for their respective tasks.
  • We achieved competitive accuracy at a fraction of the computational cost. On benchmarks like Split CIFAR-10, our LeNet-based model demonstrates that a neuroscience-aligned architecture can be both highly effective and computationally efficient in a continual learning setting.
Related Work
Continual learning (CL) aims to develop systems that learn from a non-stationary data stream without catastrophic forgetting. In this section, we review key literature spanning computational neuroscience and machine learning, focusing on complementary learning systems, hippocampal memory motifs, and modern CL strategies.
Complementary Learning Systems (CLS).
The CLS theory posits a dual-memory system: a fast hippocampal module for episodic learning and a slow neocortical system for generalization [1, 10]. This foundational concept has influenced many biologically inspired CL frameworks. Examples include hybrid cortico-hippocampal networks that combine feedforward and spiking layers [11], DualNet’s distinct fast/slow learners [12], FearNet’s use of generative replay [13], and Triple Memory Networks that model hippocampus-PFC interactions with replay and consolidation [14].
Hippocampal Replay.
Replay of past experiences is a powerful anti-forgetting mechanism, mirroring the brain’s hippocampal replay during sleep and wakefulness [15]. In machine learning, this is realized through experience replay, which explicitly rehearses stored examples [16], or generative replay, which trains a network to synthesize pseudo-data [17]. We adopt prioritized experience replay, inspired by findings that emphasizing important memories enhances learning efficiency in both reinforcement learning [18] and biological systems [19]. The ability to generate contrastive data during replay is considered crucial for preventing representational collapse [20].
Modern CL Architectures and Strategies.
Many CL frameworks are categorized as parameter isolation, rehearsal-based, or regularization-based [21].
  • Regularization methods add a loss penalty to prevent changes to weights important for past tasks, as seen in Elastic Weight Consolidation (EWC) [22], Learning without Forgetting (LwF) [23], and Synaptic Intelligence (SI) [24].
  • Parameter isolation methods assign distinct weights to different tasks, using techniques like network expansion in Progressive Nets [25] or hard attention masks [26, 27, 28].
  • Rehearsal methods use a buffer of past data [29, 30, 31] or a generative model [17, 32, 33]. Hybrid models like Progress & Compress [34] and CLEAR [30] combine these strategies effectively. Our approach synergizes these ideas, using architectural modularity, sparsity, and replay.
Mixture-of-Experts (MoE) and Gating.
MoE models enable modular specialization by routing inputs to specialized subnetworks [35, 36], a principle recently applied to CL [37, 38] and large-scale models [39, 40]. Stability is often improved via adaptive routing, for instance by annealing a routing temperature [41, 42]. Our DG-Gated MoE is distinct in that expert selection emerges via cosine similarity to prototypes, bypassing an explicit learned gating network.
Prototype-Based Replay.
Storing class prototypes instead of raw data is a memory-efficient rehearsal strategy pioneered by iCaRL [29]. Subsequent work has combined prototype anchoring with contrastive objectives to preserve the geometry of the feature space across tasks [43, 44]. We follow this line, maintaining online exponential moving average (EMA) prototypes and applying a push-pull contrastive loss during replay.
These prior threads motivate a more tightly coupled design: rather than treating biologically inspired modules as loose metaphors, we embed them directly into representation, routing, and training dynamics in a unified hippocampal-inspired dual-memory MoE.
Biologically Inspired Motifs and Dual-Memory Training
Our design is grounded in the hippocampal trisynaptic circuit (DG $\rightarrow$ CA3 $\rightarrow$ CA1) and the theory of complementary learning systems (CLS) [1, 10]. We operationalize these motifs as inductive scaffolds: structured encoding, sparse separation, nonlinear refinement/completion, and integration, which together feed into prototype-based routing and consolidation. Replay-inspired mechanisms and prototype anchoring provide retention and interference mitigation, while modular gating arises naturally from the quasi-orthogonal codes produced by DG-like separation. Learning is organized as a two-phase dual-memory schedule—specialization followed by consolidation—whose detailed losses are defined in the Architecture section.
Dual-Memory Realization. We operationalize the complementary learning system as a two-phase training schedule: Phase I rapidly specializes, forming and updating task prototypes and shaping DG gating, while Phase II consolidates prior knowledge via contrastive alignment, mitigating interference and preventing degradation.
Grid Cell Encoding
Grid-cell-like representations from the entorhinal cortex provide a structured relational prior that has been shown to enhance spatial and contextual reasoning in downstream tasks [45, 46]. We implement this by applying $M=4$ parallel $1\times 1$ convolutions with learned phase offsets to backbone features $\mathbf{f}$:
$\mathbf{g}_{m}=\sin(\mathbf{W}_{m}\,\mathbf{f}+\phi_{m}),\quad m=1,\dots,4,\qquad\mathbf{g}=[\mathbf{g}_{1};\dots;\mathbf{g}_{4}].$
This hexagonal-style embedding enhances discrimination and imposes structured relational priors before gating, similar in spirit to positional priors used in other continual and representation learning works.
Dentate Gyrus: Sparse Pattern Separation
To reduce interference between tasks, we draw on the role of the dentate gyrus in orthogonalizing inputs via sparse coding [4, 47]. The encoded representation $\mathbf{g}$ is projected and sparsified:
$\mathbf{z}$ $=\mathrm{ReLU}(\mathbf{W}_{\mathrm{DG}}\mathbf{g}+\mathbf{b}_{\mathrm{DG}}),$ (1)
$\mathbf{p}_{\mathrm{sep}}$ $=\mathrm{TopK}(\mathrm{LayerNorm}(\mathbf{z}),\,k),\quad k=\left\lfloor 0.05\,\dim(\mathbf{z})\right\rfloor.$
This top-$k$ mechanism approximates inhibitory microcircuitry to produce quasi-orthogonal sparse codes, creating natural modularization and enabling routing without explicit task labels, akin to capacity partitioning in MoE and gating literature [35, 36].
CA3: Nonlinear Refinement / Completion
The sparse separation codes are refined through a feed-forward MLP to approximate the pattern completion function attributed to CA3 [48]. Unlike classical recurrent attractor implementations [49], we use:
$\mathbf{p}_{\mathrm{comp}}=\mathrm{LayerNorm}\bigl{(}\mathrm{ReLU}(\mathbf{W}_{2}\,\mathrm{ReLU}(\mathbf{W}_{1}\,\mathbf{p}_{\mathrm{sep}}+\mathbf{b}_{1})+\mathbf{b}_{2})\bigr{)},$
providing nonlinear task-specific refinement while maintaining stability in the face of non-i.i.d. data [50, 51].
CA1: Integration
Separation and completion are fused to form a consolidated representation:
$\mathbf{u}=[\mathbf{p}_{\mathrm{sep}};\mathbf{p}_{\mathrm{comp}}],$
which serves as the basis for both prediction and prototype anchoring. This mirrors CA1’s role in integrating pattern-separated and completed signals before driving downstream systems.
Prototype Maintenance and Replay Motivation
Each expert maintains a prototype $\mathbf{u}_{i}$ updated during specialization via exponential moving average of its DG separation code:
$\mathbf{u}_{i}\leftarrow(1-\mu)\,\mathbf{u}_{i}+\mu\,\mathbf{p}_{\mathrm{sep}}^{(i)},\quad\mu=0.01.$
This echoes prototype-based rehearsal approaches [29, 43] and supplies stable references for both gating and later consolidation. The use of replay-like alignment and contrastive objectives (detailed in Section The Architecture) is inspired by hippocampal replay phenomena that reinforce important memories during offline phases [15, 18, 19].
From Motifs to Mechanism.
The preceding motifs define the representations and anchors—grid-cell priors, sparse pattern-separated codes, refined completions, and maintained prototypes—that the model builds on. The next section gives the concrete architectural instantiation, including how these pieces are composed into a DG-Gated Mixture-of-Experts and how the two-phase dual-memory objectives are applied.
The Architecture
Our system is a DG-Gated Mixture-of-Experts (MoE) designed to mitigate catastrophic forgetting by partitioning capacity into $N$ specialized experts. Each expert implements a differentiable analogue of the hippocampal trisynaptic loop (DG $\rightarrow$ CA3 $\rightarrow$ CA1), combining sparse separation, nonlinear refinement, and integration. Expert selection is driven by similarity between current DG codes and per-expert prototypes, enabling dynamic specialization without an explicit learned gating network.
Feature Extraction
A lightweight CNN backbone (e.g., a LeNet) maps each input to feature maps $\mathbf{f}$. The grid-cell encoding, sparse DG separation, refinement, and integration are as described in Section Biologically Inspired Motifs and Dual-Memory Training.
Gating & Routing
Each expert’s prototype $\mathbf{u}_{i}$ is updated during Phase I via EMA of its DG separation code:
$\mathbf{u}_{i}\leftarrow(1-\mu)\,\mathbf{u}_{i}+\mu\,\mathbf{p}_{\mathrm{sep}}^{(i)},\quad\mu=0.01.$
At inference, similarity scores
$s_{i}=\cos(\mathbf{p}_{\mathrm{sep}}^{(i)},\mathbf{u}_{i})$
are converted to gating weights via:
  • Soft gating: $\alpha_{i}\propto\exp(s_{i}/\tau)$.
  • Hard gating: $\alpha_{i}=\mathbf{1}\{i=\arg\max_{j}s_{j}\}$.
  • Top-2: select top two and renormalize.
  • Hybrid: softmax over top-$k$ similarities.
Outputs from experts are aggregated according to $\{\alpha_{i}\}$ to form the final prediction.
Training Objectives
Training proceeds in two phases.
Phase I: Specialized Encoding
The classification loss is augmented with auxiliary terms. The prototype proximity loss is denoted $\mathcal{L}_{\mathrm{intra}}$ to reflect push–pull dynamics:
$\mathcal{L}_{\mathrm{intra}}$ $=\sum_{i,j}\mathbf{1}_{[y_{i}=y_{j}]}\|\mathbf{p}_{i}-\mathbf{p}_{j}\|^{2}$ (2)
$\quad-\lambda\sum_{i,j}\mathbf{1}_{[y_{i}\neq y_{j}]}\bigl{[}\max\{0,\,m-\|\mathbf{p}_{i}-\mathbf{p}_{j}\|\}\bigr{]}^{2}.$ (3)
Additional losses are:
  • $\mathcal{L}_{\mathrm{replay}}$: cross-entropy on replayed examples.
  • $\mathcal{L}_{\mathrm{distill}}$: feature distillation aligning to frozen snapshots.
  • $\mathcal{L}_{\mathrm{EWC}}$: Elastic Weight Consolidation regularizer.
  • $\mathcal{L}_{\mathrm{sparsity}}$: encourages target DG activation sparsity.
Thus:
$\mathcal{L}_{\mathrm{phase1}}$ $=\mathcal{L}_{\mathrm{cls}}+\alpha_{\mathrm{intra}}\,\mathcal{L}_{\mathrm{intra}}+\alpha_{\mathrm{rep}}\,\mathcal{L}_{\mathrm{replay}}$
$\quad+\alpha_{\mathrm{dist}}\,\mathcal{L}_{\mathrm{distill}}+\alpha_{\mathrm{ewc}}\,\mathcal{L}_{\mathrm{EWC}}+\alpha_{s}\,\mathcal{L}_{\mathrm{sparsity}}.$ (4)
Phase II: Consolidation & Contrastive Prototypes
After freezing non-DG parameters, a prototype-based contrastive loss reinforces current-task alignment while suppressing interference:
$\mathcal{L}_{\mathrm{contrastive}}$ $=\bigl{(}1-\cos(\mathbf{p}_{\mathrm{sep}}^{(t)},\mathbf{u}_{t})\bigr{)}$
$\quad+\sum_{j\neq t}\bigl{[}\max\{0,\;\cos(\mathbf{p}_{\mathrm{sep}}^{(j)},\mathbf{u}_{j})-m\}\bigr{]}.$ (5)
$\mathcal{L}_{\mathrm{phase2}}=\alpha_{\mathrm{contrastive}}\,\mathcal{L}_{\mathrm{contrastive}}.$
Full Objective
The overall loss combines both phases and retention mechanisms:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{\mathrm{phase1}}+\lambda_{2}\mathcal{L}_{\mathrm{phase2}}+\lambda_{3}\mathcal{L}_{\mathrm{EWC}}+\lambda_{4}\mathcal{L}_{\mathrm{replay}}.$
Results
TABLE I: Comparison of accuracy (%), and FLOPs (M) on Split CIFAR-10 and Split TinyImageNet under different memory and compute settings. The baselines have been updated based on new results for Split CIFAR-10.
Method Buffer Sparsity Split CIFAR-10 Split TinyImageNet
Size Task-IL ↑ Class-IL ↑ FLOPs (M) ↓ Task-IL ↑ Class-IL ↑ FLOPs (M) ↓
Upper Bound
Joint $97.18\pm 0.12$ $87.00\pm 1.03$ 219.60 $82.04\pm 0.3$ $60.21\pm 0.3$ 850.72
Regularization-Based Methods
SGD (Fine-tuning) $66.08\pm 7.30$ $19.53\pm 0.07$ 235.66 $4.79\pm 0.43$ $0.32\pm 0.02$ 866.77
LwF $63.89\pm 1.54$ $19.40\pm 0.15$ 235.66 $21.43\pm 0.28$ $8.54\pm 0.20$ 866.77
SI $62.16\pm 4.91$ $19.29\pm 0.18$ 219.60 $21.40\pm 0.19$ $8.56\pm 0.08$ 850.72
Replay-Based Methods
A-GEM 200 $85.35\pm 2.54$ $19.66\pm 0.19$ 235.66 $22.31\pm 0.94$ $8.89\pm 0.10$ 866.77
500 $85.89\pm 2.12$ $19.83\pm 0.46$ 235.66 $23.17\pm 0.49$ $8.70\pm 0.08$ 866.77
ER 200 $90.07\pm 0.83$ $47.77\pm 1.69$ 235.66 $31.20\pm 0.10$ $8.44\pm 0.25$ 866.77
500 $92.62\pm 0.28$ $61.48\pm 3.54$ 235.66 $31.74\pm 0.29$ $8.62\pm 0.15$ 866.77
DER++ 200 $89.25\pm 0.52$ $56.73\pm 1.46$ 235.66 $33.10\pm 0.94$ $8.82\pm 0.20$ 866.77
500 $91.21\pm 0.25$ $66.49\pm 1.53$ 235.66 $32.74\pm 0.92$ $8.78\pm 0.07$ 866.77
Architectural & Hybrid Methods
iCaRL 200 $94.03\pm 0.46$ $47.54\pm 3.28$ 219.60 $24.00\pm 0.37$ $8.87\pm 0.12$ 850.72
500 $94.36\pm 0.26$ $48.34\pm 1.49$ 219.60 $23.25\pm 0.91$ $8.71\pm 0.05$ 850.72
FDR 200 $91.23\pm 0.50$ $29.51\pm 1.95$ 236.36 $35.90\pm 0.33$ $8.38\pm 0.10$ 867.48
500 $92.34\pm 1.26$ $28.73\pm 2.82$ 236.36 $41.37\pm 0.50$ $8.56\pm 0.12$ 867.48
HAL 200 $77.97\pm 2.42$ $34.11\pm 5.67$ 219.60 $21.92\pm 0.57$ $8.52\pm 0.18$ 850.72
500 $84.55\pm 1.32$ $41.98\pm 3.01$ 219.60 $21.49\pm 0.62$ $8.45\pm 0.20$ 850.72
E2Net 200 $93.11\pm 0.46$ $67.23\pm 1.57$ 219.60 $40.94\pm 0.43$ $15.35\pm 0.08$ 850.72
500 $93.48\pm 0.37$ $72.19\pm 0.82$ 219.60 $45.51\pm 0.10$ $22.47\pm 0.15$ 850.72
SparCL 200 0.25 $85.63\pm 0.62$ $45.21\pm 1.10$ 37.22 $39.14\pm 0.85$ $8.98\pm 0.38$ 149.40
500 0.25 $87.38\pm 0.46$ $50.45\pm 1.22$ 37.22 $50.83\pm 0.69$ $10.48\pm 0.29$ 149.40
Our Method
HiCL Small 200 0.15 $90.24\pm 0.84$ $57.65\pm 0.49$ 38.21 $59.16\pm 0.37$ $23.28\pm 0.46$ 152.9
500 0.15 $90.36\pm 0.54$ $60.68\pm 1.05$ 38.21 $59.63\pm 0.85$ $23.51\pm 0.91$ 152.9
HiCL Large 200 0.15 $91.74\pm 0.81$ $60.48\pm 0.15$ 84.18 $\textbf{62.76}\pm 0.26$ $25.45\pm 0.14$ 336.71
500 0.15 $92.38\pm 0.49$ $65.58\pm 0.45$ 84.18 $62.60\pm 0.79$ $\textbf{25.71}\pm 0.64$ 336.71
Benchmarks and Tasks
We evaluate our model, HiCL, on two standard image-classification benchmarks, each split into a sequence of distinct tasks to simulate a continual learning environment.
  • Split CIFAR-10 [52]: The 10 classes are divided into 5 sequential tasks, with 2 new classes introduced in each task.
  • Split Tiny-ImageNet [53]: A more complex benchmark where 200 classes are split into 10 tasks of 20 classes each.
For all experiments, we follow the task-incremental learning protocol, where the model learns tasks sequentially. We use a ResNet backbone and a fixed-size memory buffer for all methods, where applicable to ensure fair and controlled comparisons.
Evaluation Metrics
Our goal is to develop a model that learns new information effectively while forgetting as little as possible, all with high computational efficiency. To measure this, we report the following metrics:
  • Accuracy (Goal: Higher is Better): We measure final average accuracy in two scenarios.
    • Task-IL Accuracy ($A_{T}$): Measures performance when the task identity is provided at test time. This evaluates the model’s ability to preserve task-specific knowledge.
    • Class-IL Accuracy ($A_{C}$): Measures performance when the task identity is unknown, forcing the model to classify among all previously seen classes. This is a more challenging and realistic measure of a model’s ability to avoid catastrophic forgetting.
  • Efficiency (Goal: Lower is Better): We report the computational cost in Mega-FLOPs (MFLOPs) for a single forward pass. This metric serves as a proxy for the model’s inference speed and complexity.
Figure 3: DG-Gated Routing is Enabled by Effective Feature Separation. (a) The Class-IL routing matrix shows that inputs are correctly routed to their corresponding expert. For example, inputs from classes 0 and 1 (Task 0) are primarily routed to Expert 0. (b) This routing is enabled by strong pattern separation in the DG layer, where the Jaccard similarity between patterns from different tasks (’Inter-Task’) is significantly lower than within the same task (’Intra-Task’).
Training Details
We use a 2‐conv + 2‐FC LeNet backbone for CIFAR‐10 and a deeper CNN for Tiny‐ImageNet. The GridCellLayer has $M=4$ sinusoidal units, DG size =1024=1024 with top‐$k$ sparsity $k=5\%$, CA3/CA1 MLP widths $(512,256)$ and $(512,256,128)$ respectively, and replay buffer of 200 samples per task. Models are trained with Adam for 10 epochs per CIFAR‐10 task (20 for Tiny‐ImageNet), $\lambda_{\mathrm{ewc}}=0.1$, and contrastive margin $m=0.2$.
Table I reports Task‑IL accuracy, Class‑IL accuracy, and inference FLOPs per forward pass for both Small and Large HiCL variants alongside a suite of baselines. We fix the buffer at 200 or 500 samples per task and evaluate on Split CIFAR‑10 (columns 4–6) and Split Tiny‑ImageNet (columns 7–9). Baselines are grouped by methodology: regularization‑based, replay‑based, architectural/hybrid, and the joint offline upper bound.
Discussion
Summary of Key Results
HiCL demonstrates that hippocampal-inspired architecture can achieve state-of-the-art continual-learning performance with high computational efficiency. A key advantage of our DG-gated design is its support for conditional computation, where the sparse DG codes select a single expert for activation per input, avoiding the need to execute all $N$ expert pipelines.
This approach drastically lowers the total computational footprint. For instance, on Split CIFAR-10, our Small model requires 10.81 MFLOPs for the shared backbone and 5.48 MFLOPs for a single expert, resulting in an efficient inference cost of only 16.3 MFLOPs. The Large model’s cost under the same conditions is 49.7 MFLOPs. On the more complex Split Tiny-ImageNet, these costs are 49.1 MFLOPs for the Small model and 172.7 MFLOPs for the Large model. While activating a single expert may offer less GPU parallelism than a batch-parallel approach, the significant reduction in total operations makes HiCL well-suited for continual learning on resource-constrained hardware.
Mechanistic Insights
The core of HiCL’s success lies in its DG‑based gating mechanism. By enforcing top‑$k$ sparsity (Equation 1) and aggregating DG outputs into per‐expert prototypes via an EMA update, each input is mapped to a near‐orthogonal code subspace [7]. At inference, cosine similarity between live codes and prototypes yields routing accuracy exceeding 95%, consistent with high‐dimensional geometry results for random sparse vectors [54]. This sparse separation ensures expert misrouting is rare and each expert specializes effectively on its assigned tasks.
The Mechanism of DG-Gated Routing
The success of our routing mechanism stems from the fact that each expert’s Dentate Gyrus (DG) module learns a unique “feature vocabulary” tailored to its specific task. For an expert trained on Task 1 (e.g., distinguishing cats from dogs), its DG module becomes highly attuned to features like fur texture, whiskers, and ear shape. In contrast, an expert trained on Task 2 (e.g., airplanes vs. ships) will learn to amplify features related to wings, fuselages, and water.
The expert’s prototype or abstraction, which is the average of all its DG codes, becomes a compact summary of this specialized vocabulary. When a new image of a cat is presented to the model, it is processed by all experts’ DG modules in parallel. However, only the DG module from the “cat/dog” expert will produce a strong and characteristic sparse activation, as the input’s features strongly resonate with its learned vocabulary. The other experts, lacking this specialization, will produce weaker and less distinct patterns.
(a) CIFAR-10 t-SNE
(b) CIFAR-10 Prototype Similarity
(c) Tiny-ImageNet t-SNE
(d) Tiny-ImageNet Prototype Similarity
Figure 4: DG Feature Separation and Prototype Orthogonality. t-SNE plots (a, c) and similarity matrices (b, d) for Split CIFAR-10 and Split Tiny-ImageNet.
This separation is quantified by the prototype similarity matrices in Figure 4b and 4d. The strong diagonal shows that each prototype is similar to itself, while the low off-diagonal values (all close to zero) confirm that the prototypes for different tasks are nearly orthogonal. Therefore, the routing decision, a simple cosine similarity check against the prototypes is not just a superficial pattern match. It is a measure of representational resonance. The input is naturally routed to the expert whose internal feature vocabulary is best suited to understand and process it, a choice made simple and reliable by the clear, quantifiable separation of the learned feature spaces.
Memory Robustness
As shown in Figure 5, HiCL degrades gracefully as the replay buffer size shrinks. For instance, reducing the buffer from 500 to 100 examples per task causes Task-IL accuracy to drop by only 2.7 percentage points (from 92.8% to 90.1%). This robustness stems from the DG’s pattern separation: each stored exemplar occupies a distinct sparse subspace, maximizing its utility during replay. Moreover, the model’s feature-distillation and EWC regularization further stabilize representations, mitigating the effect of limited memory. These results suggest that efficient episodic encoding—rather than large buffer capacity—is a key driver of continual-learning performance.
Figure 5: Accuracy vs. Memory Size on Split CIFAR-10. The plot shows the final Task-IL Accuracy (top band), Class-IL Accuracy (blue line), and the model’s internal routing accuracy (orange dashed line) as the number of stored examples per task (memory size) varies. HiCL maintains high Task-IL accuracy and shows graceful degradation in Class-IL accuracy even with smaller buffer sizes.
Limitations and Failure Modes
Despite its strengths, HiCL has several limitations. First, it operates under a Task IL assumption, requiring known task boundaries and labels during training. While Class IL inference is label free, Phase 1 still depends on task identities; fully unsupervised gating remains an open direction. Second, we use $\rho=0.05$ and $N=5/10$ experts across all tasks. . These hyperparameters may not generalize well to domains with different levels of task heterogeneity or complexity. Third, routing can be sensitive in rare cases, with less than five percent of samples showing class confusion due to overlapping DG codes. This could potentially be addressed with adaptive gating thresholds or multi prototype routing. Lastly, our experiments are limited to vision datasets with LeNet style backbones. The model’s performance on larger scale datasets or nonvisual domains such as language or reinforcement learning remains unexplored.
Future Directions
HiCL opens up several exciting paths for future research. A key direction is enabling unsupervised task discovery by leveraging clustering or contrastive pretraining on DG codes to infer task identities, moving toward fully label free Class IL. Another promising avenue is learning adaptive sparsity, where the sparsity ratio $\rho$ is dynamically adjusted per input or expert to balance representation separation and capacity as needed. For prototype maintenance, replacing exponential moving averages with more biologically grounded Hebbian updates or even transformer based attention could improve both interpretability and performance. Scaling HiCL to deeper architectures such as ResNet or Vision Transformers, and evaluating on larger benchmarks like ImageNet Split, will be essential to assess its effectiveness beyond small scale vision tasks. Finally, translating DG gating and expert routing into neuromorphic implementations using spiking neurons could enable real time, low power continual learning on edge hardware.
Broader Impact
HiCL’s demonstration that sparse hippocampal motifs can yield efficient, robust continual learners opens pathways for real‐world AI systems that must learn adaptively under resource constraints—such as autonomous robots, edge computing devices, and lifelong learning agents. By bridging neuroscience and machine learning, we hope to inspire architectures that are not only performant but also interpretable, modular, and biologically grounded.